# Generalizing fusion rules by shuffle: Symmetry-based classifications of nonlocal systems constructed from similarity transformations

### v1

@@ Abstract
Similarity transformation is a fundamental transformations, which enables one to obtain the Hermitian systems from pseudo-Hermitian systems by the invertible transformations[1, 2, 3]. This also implements a dual relationship between the nonunitary conformal field theories (CFTs) and the nonunitary CFTs describing the corresponding pseudo-Hermitian or Hermitian lattice models respectively[4, 5]. One can see a similar (or the same) dual relationship in [6, 7, 8] related to the 3d3d3d-3d correspondence[9, 10, 11]. Related to the property of the effective vacuum or ground state, the Zamolodchikov $c$-theorem in unitary CFTs[12] has been generalized to $c_{\text{eff}}$-theorem[13] with a connection to the $PT$ symmetry in non-Hermitian systems[14, 15].
One can notice the nontrivial relationship between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems in the older studies on the realizations of chiral fermions. The lattice realization of the chiral fermions is restricted by the Nielsen-Ninomiya theorem[16, 17], and the two possible ways to realize chiral fermions are breaking locality or Hermiticity. There also exists a lattice analogue in $1+1$-dimensions[18]. We note a recent review on chiral fermion[19], those on non-Hermitian systems[20, 21, 22] as references.
We also note the pioneering nonlocal model, Haldane-Shastry model[23, 24] and several recent research directions on non-local systems, numerical studies [25, 26, 27], long-range interacting integrable models [28, 29], extension of Lieb-Schultz-Mattis theorem[30, 31, 32, 33, 34, 35], and quantum information scrambling[36, 37, 38, 39, 40, 41, 42, 43, 44]. These studies on non-local system has been developed in distinct research fields, with different techniques and languages. Hence, one may require some general method applicable to these models, such as symmetry analysis.
In this paper, we construct and study a particular type of conserved charges, symmetry topological field theories (SymTFTs)[45], of the nonlocal Hermitian systems derived from pseudo-Hermitian systems by the similarity transformations. For readers interested in SymTFTs and generalized symmetries, we note several earlier works[46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58], recent reviews[59, 60, 61, 62] and two works[63, 64] by the authors and a collaborator. The fundamental algebraic structures are a generalization of the Verlinde formula [65, 66, 67], Eq. (16), and the corresponding SymTFTs and their modules are written as Eq. (22) and Eq. (30). Their implications to RGs are summarized in Fig. 1. However, the defect or boundary phenomena are more subtle because of the nonlocality, as we demonstrate in Sec. V.
@@ II Nonunitary CFTs and their symmetry
where $\{|\alpha,M\rangle\}$ is the right eigenstates of the Virasoro generator $L_{0}(\in\{L_{m}\}_{m\in\mathbb{Z}})$ producing the corresponding Verma module labelled by the index $M$ of descendant fields, and the left eigenstates $\{\tilde{\langle\alpha|}\}$ is its linear dual satisfying the relation $\widetilde{\langle\alpha,m|}|\beta,n\rangle=\delta_{\alpha,\beta}\delta_{m,n}$. We also note the form of the Virasoro algebra
@@
where $c$ is the central charge and $\delta_{a,b}$ is the Kronekker delta. One can also implement the chiral projection operators $P_{\alpha,m}=|\alpha,m\rangle\widetilde{\langle\alpha,m|}$. In a similar way, one can also implement their antichiral analogue $\overline{P_{\alpha,\overline{M}}}$. Hence, we introduce the following idempotents,
@@
where $I$ is the identity operator corresponding to the vacuum and $N$ is the so-called nonnegative integer matrix, specifying the fusion rule $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. It should be stressed that even in nonunitary models, the matrix $N$ takes nonnegative integer values. In this setting, one can introduce the Verlinde line operator as follows[48],
@@
where these operators also satisfy the relation $\mathcal{Q}_{\alpha}\times\mathcal{Q}_{\beta}=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\mathcal{Q}_{\gamma}$ by intepreting the symbol $\times$ as multiplication of operators. It should be remarked that $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}_{\alpha}$ is as a set of linear operator over the complex number field $\mathbb{C}$ following the pioneering work[48]. By definition, these operators $\{\mathcal{Q}_{\alpha}\}$ defined by the projections commutes with the CFT Hamiltonian, and one can interprete them as conseved charges or SymTFTs. For latter use, we also note a relation,
@@
where $\{|\alpha\rangle\rangle\}$ is the Ishibashi states[77] satisfying the following relation,
$\widetilde{\langle\langle\alpha|}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}.$ (13)
By applying the Verlinde formula, the symmetry operators act on the module in the following way[78],
@@
This matrix is less familar, but it has already appeared in [65, 66], and the section $8.19.$ of [67], and it has been known to be integers. Moreover, we demonstrate that there exists the corresponding algebra, $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, which is ring isomorphic to $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}$. Hence, if there exists a unitary CFT which has the same character as a nonunitary CFT, one can obtain the dual relation between their symmetries. Fortunately, there exists at least two series of correspondence dual under the similarity, the correspondence between symplectic fermions and Dirac fermions[4] (or related $Sp(2n)_{k}\times Sp(2k)_{n}=SO(4nk)_{1}$ duality[81, 82]), and those between $M(2,2k+3)$ minimal models and $Osp(1,2)_{k}$ models (or related $W$-minimal models and $Osp(1,2n)_{k}$ models more generally[8]). We expect the same type of dualities for the models with positive $\{S_{o,\alpha}\}_{\alpha}$, and we write the corresponding SymTFTs and their modules.
@@ III Characters and Symmetry operators by shuffle
We have assumed the Hermiticity for the quantum states, $\langle\alpha,M|^{[o]}=\left(|\alpha,M\rangle^{[o]}\right)^{\dagger}$, whereas they are defined by vector and linear dual in nonunitary CFTs. This is highly similar to the pseudo-Hermiticity [21]. By applying the same procedure for the antichiral parts, we introduce the following projection operators,
@@
This expression induces the ring isomorphism $\{P_{\alpha}\}=\{P_{\alpha}^{[j]}\}$.
By definition, this shuffle operation induces the following symmetry operators,
@@
by applying the definition of $N^{[o]}$, a generalization of Verlinde formula. Hence, we have demonstrated the implication of the fusion coefficient defined by the effective vacuum, only by assuming the unitarity. We also note that $\mathcal{Q}_{o}^{[o]}=\sum_{o}P_{\beta}^{[o]}$ behaves as the identity in $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$ and the roles of $I$ and $o$ are shuffled in this sense. In the next section, we demonstrate the naturalness of the algebra $\mathbf{A}^{[o]}$, by identifying their quantum dimensions and modules (or smeared BCFTs) and by studying their implications to RGs.
It should be stressed that the condition $\{S_{j,\beta}\neq 0\}_{\beta}$ implies the invertibility of $\mathcal{Q}_{j}$ in linear algebra. It is well-known that the invertibility of the symmetry plays a significant role in studying gauging operations[57, 58], and our work provides a generalization respecting the linear algebraic structure of quantum systems. Moreover, one can obtain the following relations,
@@ III.1 $M(2,5)$ minimal model and $Osp(1,2)_{1}$ WZW model
where $\mathcal{Q}_{o}^{[o]}$ is the identity operator. As can be seen from the above expression, the roles of $I$ and $o$ are shuffled and their conformal dimensions are $h_{o}^{[o]}=0$, $h_{I}^{[o]}=1/5$. These conformal dimensions and the effective central charge $c_{\text{eff}}=2/5$ match with those of $Osp(1,2)_{1}$ Wess-Zumino-Witten (WZW) model[7, 8]. Hence, the above $\mathbf{A}^{[o]}$ provides a natural expression of SymTFT of the $Osp(1,2)_{1}$ WZW model.
@@ IV Ring isomorphism and application to RGs
In this section, we study the consequences of the similarity transformation in SymTFTs combined with massive and massless RG. We note the ring isomorphism as $\iota:\mathbf{A}\rightarrow\mathbf{A}^{[o]}$ in the ultraviolet(UV) theories. At the algebraic level, the massless RG corresponds to a ring homomorphism (or functor), and this aspect has been studied in the name RG domain wall[97, 55, 98] or gapped domain wall[99, 100, 101, 102, 103]. On the other hand, the massive RG corresponds to the operation taking module (or smeared BCFT) of the unbroken symmetry subring of the original UV theory[104, 105, 106, 107, 72, 108, 73, 74, 75, 76, 109, 110, 63, 64].
@@ IV.1 Massless RG flow
In this subsection, we note the implications of the similarity transformations realized as ring isomorphism. We denotes an infrared (IR) theories as $\mathbf{A^{\prime}}$ and $\mathbf{A^{\prime}}^{[o]}$ and their similarity transformation as $\iota^{\prime}:\mathbf{A}^{\prime}\rightarrow\mathbf{A^{\prime}}^{[o]}$.
First, we assume the massless RG or ring homomorphism between a UV theory $\mathbf{A}$ to an IR theory $\mathbf{A^{\prime}}$[100, 111, 103]. In this setting, one can implement the ring homomorphism of the corresponding nonlocal unitary model as
@@
Interestingly, the simple objects in $\mathbf{A}_{\text{ub}}$ are not mapped to simple objects in $\iota(\mathbf{A}_{\text{ub}})$ in general. By combining this observation with the homomorphism $\rho$, one can notice that the nonsimple (or nonlocal) objects in $\mathbf{A}$ can be simple objects in the IR theory $\mathbf{A^{\prime}}_{\text{ub}}^{[o]}$. At the level of linear algebra or ring theory, this phenomenon is understandable, but its phenomenologies or categorical descriptions are less clear.
@@ IV.2 Smeared BCFT and massive RG flow
Hence, by assuming the massive RG is realized as reduction of the symmetry from $\mathbf{A}^{[o]}$ to $\mathbf{A}_{\text{ub}}^{[o]}$[72, 73, 74, 75, 76, 63, 64], one can classify the gapped phase as $\mathbf{A}_{\text{ub}}^{[o]}$ invariant subspace of the Hilbert space spaned by $\{|\alpha\rangle^{[o]}\}$.
In this section, the unitarity combined with recent developments in BCFT[78, 72] has played a fundamental role, and the existing arguments in the smeared BCFT is applicable to our cases. However, the Cardy’s condition which determines the boundary critical phenomena is broken, because the amplitude can become outside of NIM-rep[80],
$\begin{split}Z_{\alpha,\beta}^{[o]}(\tau)&=\langle\alpha|^{[o]}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c_{\text{eff}}/12)}|\beta\rangle^{[o]}\\ &=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}\chi_{\gamma}(-1/\tau)\end{split}$ (32)
@@ V Difficulty in boundary or defect phenomena
As we have observed in the previous section, the ring $\mathbf{A}^{[o]}$ outside of the NIM-rep[79, 80] cannot describe the defect or boundary phenomena straightforwardly. However, there still exist similarity transformations for the non-Hermitian system with such objects. For simplicity, we concentrate our attention on the annulus nonunitary CFT determined by $Z_{\alpha,\beta}$.
In this setting, by the open-closed duality[80], one can map the Hilbert space to that of the chiral CFT, $\oplus_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. Hence, we propose a conjecture:
[…]
Under this assumption, one can apply the existing discussions on the corresponding vertex-operator algebraic arguments to the unexplored boundary physics of the nonlocal models. More technically, the doubling trick[113, 114] will play a fundamental role in studying the correlation functions and related quantities. We also point out that the above similarity transformation induced by some invertible operator can be regarded as a generalization of the large gauge transformation induced by unitary operators, known as Lieb-Schultz-Mattis operators[115, 116, 117, 118] [^1] We thank Masaki Oshikawa for the corresponding discussion.. This interpretation has appeared in [120].
By choosing the left and right boundary conditions carefully for $Z_{\alpha,\beta}^{[o]}$, one can also avoid the negative amplitude. This nontrivial (or nonlocal) relation between the nonnegative amplitude condition and the boundary conditions can be interpreted as a signal of the nonlocality. However, to our knowledge, the relationship between this nonlocality and the similarity transformation has not been studied. For example, in $M(2,5)$ minimal model, one can obtain the same characters from the pairs $(Z_{o,o}^{[o]},Z_{I,o})$, $(Z_{o,I}^{[o]},Z_{I,I})$, but it is impossible to reproduce $Z_{o,o}$ from $Z^{[o]}_{\alpha,\beta}$.
The phenomenon seems to be related to the boundary sensitity[121, 122, 123] or quantum skin effect[124] in non-Hermitian systems, but their precise relationship is still unclear. We also note some works studying similar problems arising from the boundary conditions of pseudo-Hermitian systems[125, 126], and a pioneering work pointing out its relation to quantum groups[127].

### v2

@@ Abstract
Similarity transformation is a fundamental transformation, which enables one to obtain the Hermitian systems from pseudo-Hermitian systems by the invertible transformations[1, 2, 3]. This also implements a dual relationship between the nonunitary conformal field theories (CFTs) and the nonunitary CFTs describing the corresponding lattice models[4, 5]. One can see a similar (or the same) dual relationship in [6, 7, 8] related to the 3d3d3d-3d correspondence[9, 10, 11]. Related to the property of the effective vacuum or ground state, the Zamolodchikov $c$-theorem in unitary CFTs[12] has been generalized to $c_{\text{eff}}$-theorem[13] with a connection to the $PT$ symmetry in non-Hermitian systems[14, 15].
One can notice the nontrivial relationship between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems in the older studies on the realizations of chiral fermions[16, 17, 18, 19, 20, 21, 22]. The lattice realization of the chiral fermions is restricted by the Nielsen-Ninomiya theorem[23, 24], and the two possible ways to realize chiral fermions are breaking locality or Hermiticity. There also exists a lattice analogue in $1+1$-dimensions[25]. We note a recent review on chiral fermion[26] and those on non-Hermitian systems[27, 28, 29] as references.
We also note the pioneering nonlocal model, Haldane-Shastry model[30, 31] and several recent research directions on non-local systems, numerical studies [32, 33, 34], long-range interacting integrable models [35, 36], extension of Lieb-Schultz-Mattis theorem[37, 38, 39, 40, 41, 42], and quantum information scrambling[43, 44, 45, 46, 47, 48, 49, 50, 51]. These studies have been developed in distinct research fields, with different techniques and languages. Hence, one may require some general method applicable to these models, such as symmetry analysis.
In this paper, we construct and study a particular type of conserved charges, symmetry topological field theories (SymTFTs)[52], of the nonlocal Hermitian systems derived from pseudo-Hermitian systems by the similarity transformations. For readers interested in SymTFTs and generalized symmetries, we note several earlier works[53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], recent reviews[66, 67, 68, 69] and two works[70, 71] by the authors and a collaborator. The fundamental algebraic structure is a generalization of the Verlinde formula [72, 73, 74], Eq. (14), and the corresponding SymTFTs and their modules are written as Eq. (20) and Eq. (28). Their implications to RGs are summarized in Fig. 1. However, the defect or boundary phenomena are more subtle because of the nonlocality, as we demonstrate in Sec. V.
@@ II Nonunitary CFTs and their symmetry
where $\{|\alpha,M\rangle\}$ is the right eigenstates of the Virasoro generator $L_{0}(\in\{L_{m}\}_{m\in\mathbb{Z}})$ producing the corresponding Verma module labelled by the index $M$ of descendant fields, and the left eigenstates $\{\tilde{\langle\alpha|}\}$ is its linear dual satisfying the relation $\widetilde{\langle\alpha,M|}|\beta,M^{\prime}\rangle=\delta_{\alpha,\beta}\delta_{M,M^{\prime}}$. We also note the form of the Virasoro algebra
@@
where $c$ is the central charge and $\delta_{a,b}$ is the Kronekker delta. One can also implement the chiral projection operators $P_{\alpha,M}=|\alpha,M\rangle\widetilde{\langle\alpha,M|}$. In a similar way, one can also implement their antichiral analogue $\overline{P_{\alpha,\overline{M}}}$. Hence, we introduce the following idempotent,
@@
where $I$ is the identity operator corresponding to the vacuum and $N$ is the so-called nonnegative integer matrix (NIM), specifying the fusion rule $\alpha\times\beta=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. We stress that even in nonunitary models, the matrix $N$ is NIM. In this setting, one can introduce the Verlinde line operator as follows[55],
@@
where these operators also satisfy the relation $\mathcal{Q}_{\alpha}\times\mathcal{Q}_{\beta}=\sum_{\gamma}N^{\gamma}_{\alpha,\beta}\mathcal{Q}_{\gamma}$ by intepreting the symbol $\times$ as multiplication of operators. Remarkablly, $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}_{\alpha}$ is a set of linear operators over the complex number field $\mathbb{C}$ following the pioneering work[55]. By definition, $\{\mathcal{Q}_{\alpha}\}$ defined by the projections commutes with the CFT Hamiltonian, and one can interprete them as conseved charges or SymTFTs. For latter use, we also note a relation,
@@
where $\{|\alpha\rangle\rangle\}$ is the Ishibashi states[84] satisfying the relation, $\widetilde{\langle\langle\alpha|}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}.$ By applying the Verlinde formula, the symmetry operators act on the module in the following way[85],
@@
This matrix is less familar, but it has already appeared in [72, 73], and the section $8.19.$ of [74], and it has been known to be integers. Moreover, we demonstrate that there exists the corresponding algebra, $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, which is ring isomorphic to $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}$. Hence, if there exists a unitary CFT that has the same character as a nonunitary CFT, one can obtain the dual relation between their symmetries. Fortunately, there exists at least two series of orrespondence dual under the similarity, the correspondence between symplectic fermions and Dirac fermions[4] (or related $Sp(2n)_{k}\times Sp(2k)_{n}=SO(4nk)_{1}$ duality[88, 89]), and those between $M(2,2k+3)$ minimal models and $Osp(1,2)_{k}$ models (or related $W$-minimal models and $Osp(1,2n)_{k}$ models[8]). We expect the same type of dualities for the models with positive $\{S_{o,\alpha}\}_{\alpha}$, and we write the corresponding SymTFTs and their modules.
@@ III Characters and Symmetry operators by shuffle
We have assumed the Hermiticity for the quantum states, $\langle\alpha,M|^{[o]}=\left(|\alpha,M\rangle^{[o]}\right)^{\dagger}$, whereas they are defined by vector and linear dual in nonunitary CFTs. By applying the same procedure for the antichiral parts, we introduce the following idempotent,
@@
This expression induces the ring isomorphism $\{P_{\alpha}\}=\{P_{\alpha}^{[o]}\}$.
By definition, this shuffle operation implements the following symmetry operators,
@@
by applying the definition of $N^{[o]}$, a generalization of Verlinde formula. Consequently, we have demonstrated the implication of the fusion coefficient defined by $o$, only by assuming the unitarity. We also note that $\mathcal{Q}_{o}^{[o]}=\sum_{o}P_{\beta}^{[o]}$ behaves as the identity in $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$ and the roles of $I$ and $o$ are shuffled. In the next section, we demonstrate the naturalness of the algebra $\mathbf{A}^{[o]}$, by identifying its quantum dimensions and modules (or smeared BCFTs) and by studying their implications to RGs.
It should be stressed that the condition $\{S_{j,\beta}\neq 0\}_{\beta}$ implies the invertibility of $\mathcal{Q}_{j}$ in linear algebra. It is well-known that the invertibility of the symmetry plays a significant role in studying gauging operations[64, 65], and our argument provides a generalization respecting the linear algebraic structure of quantum systems. Moreover, one can obtain the following relations,
@@ III.1 $M(2,5)$ minimal model and $Osp(1,2)_{1}$ WZW model
where $\mathcal{Q}_{o}^{[o]}$ is the identity operator. As can be seen from the above expression, the roles of $I$ and $o$ are shuffled, and their conformal dimensions are $h_{o}^{[o]}=0$, $h_{I}^{[o]}=1/5$. These conformal dimensions and the effective central charge $c_{\text{eff}}=2/5$ match with those of $Osp(1,2)_{1}$ Wess-Zumino-Witten (WZW) model[7, 8]. Hence, the above $\mathbf{A}^{[o]}$ provides a natural expression of SymTFT of the $Osp(1,2)_{1}$ WZW model.
@@ IV Ring isomorphism and application to RGs
In this section, we study the consequences of the similarity transformation in SymTFTs combined with massive and massless RG. We denote the ring isomorphism as $\iota:\mathbf{A}\rightarrow\mathbf{A}^{[o]}$ in the ultraviolet(UV) theories. At the algebraic level, the massless RG corresponds to a ring homomorphism (or functor), and this aspect has been studied in the name RG domain wall[104, 62, 105] or gapped domain wall[106, 107, 108, 109, 110]. On the other hand, the massive RG corresponds to the operation taking module of the unbroken symmetry subring of the original UV theory[111, 112, 113, 114, 79, 115, 80, 81, 82, 83, 116, 117, 70, 71].
@@ IV.1 Massless RG flow
In this subsection, we note the implications of the similarity transformations realized as ring isomorphism. We denote infrared (IR) theories as $\mathbf{A^{\prime}}$ and $\mathbf{A^{\prime}}^{[o]}$ and their similarity transformation as $\iota^{\prime}:\mathbf{A}^{\prime}\rightarrow\mathbf{A^{\prime}}^{[o]}$.
First, we assume the massless RG or ring homomorphism $\rho$, $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$[107, 118, 110]. In this setting, one can implement the ring homomorphism of the corresponding nonlocal unitary model as
@@
Interestingly, the simple objects in $\mathbf{A}_{\text{ub}}$ are not mapped to simple objects in $\iota(\mathbf{A}_{\text{ub}})$ in general. By combining this observation with the homomorphism $\rho$, one can notice that the nonsimple (or nonlocal) objects in $\mathbf{A}$ can be simple objects in the IR theory $\mathbf{A^{\prime}}_{\text{ub}}^{[o]}$. At the level of linear algebra, this phenomenon is understandable, but its phenomenologies or categorical descriptions are less clear.
@@ IV.2 Smeared BCFT and massive RG flow
where we call the right-hand side as the Graham-Watts (GW) states. Hence, by assuming the massive RG is realized as reduction from $\mathbf{A}^{[o]}$ to $\mathbf{A}_{\text{ub}}^{[o]}$[79, 80, 81, 82, 83, 70, 71], one can classify the gapped phase as $\mathbf{A}_{\text{ub}}^{[o]}$ invariant subspace of the Hilbert space spaned by $\{|\alpha\rangle^{[o]}\}$. In this section, the unitarity combined with recent developments in BCFT[85, 79] has played a fundamental role, and the existing arguments in the smeared BCFT is applicable to our cases. However, the Cardy’s condition which determines the boundary critical phenomena is broken, because the amplitude can become outside of NIM-rep[87],
$\begin{split}Z_{\alpha,\beta}^{[o]}(\tau)&=\langle\alpha|^{[o]}e^{\pi i\tau(L_{0}^{[o]}+\overline{L_{0}}^{[o]}-c_{\text{eff}}/12)}|\beta\rangle^{[o]}\\ &=\sum_{\gamma}N^{\gamma,[o]}_{\alpha,\beta}\chi_{\gamma}(-1/\tau)\end{split}$ (30)
@@ V Difficulty in boundary or defect phenomena
As we have observed in the previous section, the ring $\mathbf{A}^{[o]}$ outside of the NIM-rep[86, 87] cannot describe the defect or boundary phenomena straightforwardly. However, there still exist similarity transformations for the non-Hermitian system with such objects. For simplicity, we concentrate our attention on the annulus nonunitary CFT determined by $Z_{\alpha,\beta}$. In this setting, by the open-closed duality[87], one can map the Hilbert space to that of the chiral CFT, $\oplus_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. Hence, we propose a conjecture for the shuffle $\eta^{[o]}$ and the similarity transformation:
$\eta^{[o]}$ implements the similarity transformation to $\oplus_{\gamma}N^{\gamma}_{\alpha,\beta}\gamma$. (31)
Under this assumption, one can apply the existing formalisms in the chiral CFTs to the unexplored boundary physics of the nonlocal models. More technically, the doubling trick[120, 121] will play a fundamental role in studying the correlation functions and related quantities. We also point out that the above similarity transformation induced by some invertible operator can be regarded as a generalization of the large gauge transformation induced by unitary operators, known as Lieb-Schultz-Mattis operators[122, 123, 124, 125] [^1] We thank Masaki Oshikawa for the corresponding discussion.. This kind of interpretation has appeared in [127].
By choosing the left and right boundary conditions carefully for $Z_{\alpha,\beta}^{[o]}$, one can also avoid the negative amplitude. This nontrivial (or nonlocal) relation between the nonnegative amplitude condition and the boundary conditions can be interpreted as a signal of the nonlocality. For example, in the $M(2,5)$ model, one can obtain the same characters from the pairs $(Z_{o,o}^{[o]},Z_{I,o})$, $(Z_{o,I}^{[o]},Z_{I,I})$, but it is impossible to reproduce $Z_{o,o}$ from $Z^{[o]}_{\alpha,\beta}$, by taking the label $\alpha,\beta$ as that of the GW states. By introducting the unusual non-GW states $|o\rangle^{[o]}+|I\rangle^{[o]}$, $Z_{o,I+o}^{[o]}$ produces $Z_{o,o}$, but this breaks the correspondence between left and right boundaries in $Z_{o,o}$.
The phenomenon seems to be related to the boundary sensitity[128, 129, 130] or quantum skin effect[131] in non-Hermitian systems, and it is worth further study. We also note some works studying similar problems arising from the boundary conditions of pseudo-Hermitian systems[132, 133], and a pioneering work pointing out its relation to quantum groups[134].